OpenStudy (anonymous):

solve. 2/3 + t = 4/5

OpenStudy (anonymous):

I got 2/15

OpenStudy (anonymous):

2/3 + t = 4/5 t = 4/5 - 2/3 t = 2/15 You're correct.